OpenStudy (anonymous):

In a class of 25 students, the average score is 84. Six students in the class each recieved a max score of 100, and 3 students each recieved a score of 60. What is the average score of the remaining students?

OpenStudy (anonymous):

84=sum/25 6(100)+3(60)+x=sum find the sum from 1st equation then use it to find x in the 2nd question

OpenStudy (anonymous):

I do not understand the first equation. sum of what?

OpenStudy (anonymous):

sum of the marks for all the students in the class average=sum of items/number of items 84=sum/25

OpenStudy (anonymous):

so 84/25?

OpenStudy (anonymous):

sum=84*25

OpenStudy (anonymous):

are we together here we multiply both sides by 25\[84=\frac{Sum}{25}\] \[Sum=84*25=2100\]

OpenStudy (anonymous):

ok 6(100)+30(60)+x=2100?

OpenStudy (anonymous):

yes so x=?

OpenStudy (anonymous):

-300?

OpenStudy (anonymous):

jk sorry hold on its 3 not 30

OpenStudy (anonymous):

x=1320

OpenStudy (anonymous):

yes my mistake good x=1320, so what is this x is it our answer or what

OpenStudy (anonymous):

no cant be

OpenStudy (anonymous):

so look this is the sum of the students who where left out -remaining how many where they remember we used only 6+3 students

OpenStudy (anonymous):

do I divide it by 16?

OpenStudy (anonymous):

good ,yes thats it,that will be their average

OpenStudy (anonymous):

ok. took me a while to get there haha i apprecite your help

OpenStudy (anonymous):

yw